OpenStudy (anonymous):

x3-6 how to find zeroes

OpenStudy (phi):

\[ x^3 - 6 =0 \] ? if so, \[ x^3 -6 + 6 = 0 + 6 \\ x^3 = 6 \\ x= \sqrt[3]{6} \]